### v1

# World In Your Hands: A Large-Scale and Open-source Ecosystem for Learning Human-centric Manipulation in the Wild

Abstract
Large-scale pre-training is fundamental for generalization in language and vision models, but data for dexterous hand manipulation remains limited in scale and diversity, hindering policy generalization. Limited scenario diversity, misaligned modalities, and insufficient benchmarking constrain current human manipulation datasets. To address these gaps, we introduce World In Your Hands (WiYH), a large-scale open-source ecosystem for human-centric manipulation learning. WiYH includes (1) the Oracle Suite, a wearable data collection kit with an auto-labeling pipeline for accurate motion capture; (2) the WiYH Dataset, featuring over 1,000 hours of multi-modal manipulation data across hundreds of skills in diverse real-world scenarios; and (3) extensive annotations and benchmarks supporting tasks from perception to action. Furthermore, experiments based on the WiYH ecosystem show that integrating WiYH’s human-centric data significantly enhances the generalization and robustness of dexterous hand policies in tabletop manipulation tasks. We believe that World In Your Hands will bring new insights into human-centric data collection and policy learning to the community.
1 Introduction
| Dataset | Hours | Clips | Wild | Skills | Cam. Calib | RGB | Tactile | Action | Depth | Mask | Instruct. | Reason |
| Ego4D [grauman2022ego4d] | 3,670 | - | ✓ | ✓ | ✗ | ✓ | ✗ | 2D | ✗ | ✗ | ✓ | ✓ |
| HOI4D [liu2022hoi4d] | 44.4 | 4k | ✓ | ✓ | ✓ | ✓ | ✗ | 2D | ✓ | ✓ | ✗ | ✗ |
| HoloAssist [wang2023holoassist] | 166 | 2.22k | ✗ | ✗ | ✗ | ✓ | ✗ | - | ✓ | ✗ | ✓ | ✓ |
| Ego-Exo4D [grauman2024ego] | 1,286 | 5.04k | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
| CaptainCook4D [peddi2024captaincook4d] | 94.5 | 384 | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ |
| HOT3D [banerjee2025hot3d] | 13.9 | 425 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✗ | ✗ | ✗ | ✗ |
| HO-Cap [wang2024ho] | - | 64 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✓ | ✓ | ✗ | ✗ |
| Ego-ViD-5M [wang2024egovid] | - | 5M | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
| EgoDex [hoque2025egodex] | 829 | 338k | ✗ | ✓ | ✗ | ✓ | ✗ | 3D | ✗ | ✓ | ✓ | ✗ |
| WiYH (Ours) | 1045 | 125.4k | ✓ | ✓ | ✓ | ✓ | ✓ | 3D | ✓ | ✓ | ✓ | ✓ |
Large-scale pre-training has been established as a cornerstone for achieving generalization in Large Language Models [brown2020language, chowdhery2023palm, achiam2023gpt, touvron2023llama, yang2025qwen3], Vision-Language Models [radford2021learning, li2022blip, alayrac2022flamingo, li2023blip, li2024llava] and Vision-Language-Action Models [brohan2022rt, zitkovich2023rt, kim2024openvla, black2410pi0, wen2024diffusion, zhou2025vision]. However, datasets for dexterous hand manipulation remain far smaller and less diverse than language datasets, limiting progress in learning robust manipulation policies.
To address the problem of data scarcity, human-centric data collection and learning have attracted growing interest [hsieh2025dexman, zhu2025learning, routray2025vipra, zhu2025let, kareer2025egomimic]. Some approaches [chao2021dexycb, kwon2021h2o, banerjee2025hot3d] capture human hand data during object manipulation using devices such as VR systems and smart glasses, while others [grauman2022ego4d, xu2021videoclip, wang2021actionclip, he2024egovm] leverage foundation models to extract human actions from unlabeled videos of human manipulation available on the internet.
Despite these advances, current human manipulation datasets still face three major limitations, as shown in Table 1 that hinder their usefulness for dexterous manipulation research: (1) Limited Scenario Diversity: Many datasets are collected in constrained laboratory environments, lacking the diversity and complexity of real-world settings. (2) Inadequate Alignment: Existing collections often lack one or more key modalities, such as the absence of fine-grained language instructions, precise 3D hand poses, or object poses, resulting in misalignment across vision, language, and action. (3) Insufficient Benchmarking: Most datasets do not support comprehensive evaluation benchmarks that span the entire pipeline of dexterous manipulation, including scene perception, task planning, and action.
To bridge this gap and investigate in-depth how human-centric data can advance dexterous manipulation, we introduce World In Your Hands (WiYH), a large-scale open-source ecosystem for learning human-centric manipulation in the wild, as shown in Figure 1. Specifically, the WiYH ecosystem comprises three core components:
The Oracle Suite and Auto-labeling Pipeline. We introduce a lightweight and flexible wearable data collection kit, the Oracle Suite, designed to capture data without impeding the operator’s natural movements. An integrated auto-labeling pipeline enables markerless self-localization and centimeter-level motion accuracy without requiring external third-person tracking.
The WiYH Dataset: A Large-Scale Multi-modal Human-centric Dataset. We present the WiYH dataset, a large-scale multi-modal resource comprising over 1,000 hours of manipulation sequences across hundreds of skills, captured in 10 diverse industrial and daily-life scenarios such as factories, supermarkets, hotels, and restaurants. The dataset includes rich sensory streams: multi-view images, camera calibration parameters, depth maps, 3D hand poses, and 6D wrist trajectories.
Diverse Annotations and Comprehensive Benchmarks. Building upon the WiYH dataset, we provide extensive annotations, including approximately 400 hours of atomic-level action instructions and 100k instances of vision-language data for task reasoning, affordance understanding, and object localization. These annotations support a comprehensive set of benchmarks—spanning visual language understanding, world model evaluation, and vision-language-action tasks—designed to assess capabilities from scene perception to embodied operation holistically.
Based on the WiYH ecosystem, we study how human-centric data empowers dexterous manipulation learning in tabletop manipulation tasks. We train models on mixtures of Oracle Suite demonstrations and other manipulation data and evaluate the task success rate across varied environments and object layouts. Our experiments demonstrate that human-centric data substantially improves the generalization and robustness of dexterous hand policies.
[Limitation]. All manipulation tasks are evaluated in real-world environments. To increase accessibility and reproducibility, we are developing high-fidelity simulation environments and plan to release them in a future version of the dataset and benchmarks.
2 World In Your Hands Ecosystem
2.1 Oracle Suite: Human-centric Data Collection Suite
To enable large-scale, high-fidelity, human-centric multimodal data acquisition, we developed Oracle Suite, a low-cost wearable data collection system designed for human-centric embodied AI research. Specifically, as shown in Figure 2, Oracle Suite comprises a hardware system for acquiring multi-view RGB, IMU-based localization, and tactile information, alongside an autolabeling module that achieves high-precision motion capture through the fusion of IR, RGB, and IMU localization.
Hardware System Overview. Oracle Suite adopts a modular architecture integrating three primary hardware components: H-HPVHive, H-Gloves, and H-Backpack. This modular design enables flexible deployment across diverse environments, without being restricted to controlled indoor labs.
The Oracle Suite introduces major optimizations in ergonomics, sensing, and reliability:
H-FPVHive: It supports chest-mounted configuration, and integrates two fisheye cameras, two pinhole cameras, and four infrared lenses. All cameras are hardware synchronized. The infrared lenses are used for better locating the relative poses of H-Gloves.
H-Gloves: Each glove contains 6 IMUs, 5 pressure sensors, and an on-board MCU. The gloves can capture the hand motions and tactile. Moreover, it also contains 3 fisheye cameras for visual tracking and observation.
H-Backpack: It provides data storage, computation, and power supply. The computation is hosted by NVIDIA Orin.
Autolabeling Pipeline. As shown in the right part of Figure 2, the Autolabeling Pipeline comprises both online and offline algorithms. It fuses multi-modal sensor data (IR + IMU + RGB) to ensure positioning accuracy in in-the-wild scenarios. Specifically, the online localization module performs lightweight coarse positioning using visual-inertial odometry (VIO) algorithms and IR detection, while the offline stage refines these coarse estimates through visual Structure-from-Motion (SfM) algorithms. The pipeline ultimately outputs centimeter-level 6D poses for both wrists.
Data Production Efficiency. The Oracle Suite pipeline follows a collection–annotation–inspection–archiving workflow for scalable data production. A single collector operating 5 hours per day generates approximately 1.8 TB of valid multimodal data. At scale, 100 collectors over 30 days yield roughly 5.4 PB of raw data (stored in cloud repositories). Efficiency is further improved through standardized capture protocols and automated cloud-based annotation.
2.2 Data Annotation Pipeline
Data collection: Practitioners wore the Oracle Suite in each scene to record natural manipulation behavior. Each recording session lasted approximately 10 minutes and included start and end event markers. The collected signals comprise tactile signals from pressure sensors, binocular fisheye images from the left and right wrist mounts, first-person binocular fisheye images, left-hand and right-hand IMU localization data, and the corresponding camera calibration parameters. Samples of data annotations are shown in Figure 4.
Data annotation: The annotation pipeline comprises three parts: subtask (atomic task) annotation, perception annotation, and vision-language annotation.
Subtask annotation: To construct a large-scale, high-quality set of atomic tasks, we designed a four-stage hybrid annotation pipeline consisting of data processing, task definition, annotation consolidation, and post-processing:
(1) Data processing: First-person fisheye videos are segmented into 30-second clips, which serve as the basic units for semantic annotation by the video-language module;
(2) Task definition (manual annotation): The video-language module parses hierarchical operational semantics from the first-person videos and hierarchically decomposes an operation sequence into high-level tasks and atomic subtask units. Each task unit contains a preset action described in natural language and a manually annotated target object.
(3) Data checking (manual review): Outputs from the video-language module are subject to iterative manual review; any incorrect annotations are reprocessed before proceeding to the consolidation stage.
(4) Post-processing: Annotated preset-action and target-object names are assembled into complete operation instructions and are further augmented using a large language model (LLM). After subtask annotation, the raw data are organized into a hierarchical format of tasks composed of multiple atomic subtasks.
Perception annotation: To provide richer perceptual information, we pre-annotated scene elements using visual foundation models and then performed manual selection and correction. The resulting perception annotations include:
(1) Segmentation masks: instance masks are generated by applying SAM to human-provided prompt points;
(2) Depth maps: scene depth is estimated from first-person binocular images using the FoundationStereo model.
Vision-language annotation: The vision-language annotations were produced through manual curation and comprise the following tasks:
(1) Spatial Referring: For each subtask, we extract a key frame near the action start that contains the operator’s object of attention and in which the target object is not occluded by the hand, body, or other objects. Based on the subtask’s target object, spatial relational terms are added to uniquely refer to a single target region in the image.
(2) Subtask Prediction: We select tasks that contain 4–5 subtasks. The next subtask following the current one is used as the correct option. Nine incorrect options are generated by combining scene object information and the task’s preset actions; annotators then select the three incorrect options that are the most logically plausible. Additionally, Qwen-2.5VL-72B was used to generate auxiliary “thinking” data for next-action prediction, which were subsequently corrected by human annotators;
(3) Completion Verification: We select tasks containing 4-5 subtasks and randomly trim the tail of the video by 20%, 25%, 30%, or 35% of frames. The task description and the truncated video are provided as input, and human annotators label whether the video depicts a completed task.
Through the above procedures, we obtain rich annotations covering scene understanding and task planning, which constitute the foundation of a unified benchmark.
2.3 Data Statistics
Figure 3 summarizes key statistics of our WiYH dataset. We first analyze the distribution of tasks across scenes. As shown in Figure 3(a), WiYH contains a diverse set of daily environments, including banquet, laundry, logistics, hotel, department, office, supermarket, industry, cleaning, and candlelight settings. The Sankey diagram illustrates the total accumulated duration of each task within its corresponding scene.
Scenes such as supermarkets and laundry include long-horizon activities with substantial execution time, while hotels and candlelight contain shorter but highly structured routines. This diversity in task–scene composition highlights the wide operational range covered by WiYH. Action-level temporal characteristics are shown in Figure 3(b). For each scene, we compute the average execution duration of fine-grained actions. The dataset exhibits substantial variation across action categories, from short, instantaneous operations such as wipe and insert, to long-horizon manipulation such as smooth out or foldlay. These temporal differences reflect the heterogeneous complexity of real-world household and industrial workflows.
Figure 3(c) reports the distribution of annotation types. WiYH includes rich multimodal and task-level supervision, consisting of RGB, depth, action, instruction, reasoning, mask, tactile, and calibration annotations. Among them, RGB and calibration frames appear most frequently, while reasoning and tactile annotations, though smaller in quantity, provide high-level semantic cues crucial for complex manipulation and planning tasks.
To better understand the manipulation space in WiYH, Figure 3(d–e) visualizes word clouds of the annotated target objects and skills. The dataset covers a broad range of manipulable objects—including clothes, linens, cartons, cups, and tableware—spanning deformable, rigid, and articulated categories. Likewise, the skill vocabulary encompasses diverse manipulation primitives such as take, place, rotate, unfold, align, push, and search, representing both low-level motor actions and higher-level task strategies. These distributions further demonstrate that WiYH captures a wide spectrum of real-world manipulation behaviors.
3 Benchmark and Applications
3.1 Human-centric Vision-Language (HVL)
In this section, we construct three evaluation tasks based on the vision-language annotations introduced in Section 2. These tasks are designed to evaluate whether existing Multimodal Large Language Models (MLLMs) can achieve a genuine understanding of human-centric manipulation tasks in highly generalized in-the-wild environments.
We define the following three evaluation tasks and test several closed-source and open-source MLLMs, including GPT-4o [hurst2024gpt], Doubao-seed [bai2024seed], and Qwen [yang2025qwen3]. Detailed experimental setups and data statistics are provided in the supplementary material. (1) Spatial Referring (SR): Given a human manipulation image and a description of the spatial relationships among multiple objects in the scene, the model is required to output the pixel coordinates of a specified region. Performance is measured by the probability that the predicted point lies within the target area. (2) Subtask Prediction (SP): Given a piece of human manipulation video, the current task description, and four candidate options for the next action, the model must select the correct subsequent action. (3) Completion Verification (CV): Given a piece of human manipulation video and the current task description, the model needs to determine whether the task has been completed.
As shown in Table 2, Qwen3-VL-Plus, which was pre-trained on spatial understanding data, achieves the best performance on the Spatial Referring task. However, scores of all MLLMs are below 0.5, indicating that current MLLMs still lack sufficient capability in interpreting fine-grained spatial relationships in manipulation scenarios. On the Completion Verification task, all models perform slightly above 0.5, suggesting that their ability to accurately judge task completion in such contexts remains limited. These experimental results collectively demonstrate that general-purpose, vanilla VLLMs still fall short in embodied tasks. Our WiYH dataset, with its rich annotations of human activities and spatial relationships, presents a promising foundation for future embodied pre-training.
| Model | SR (%) | SP (%) | CV (%) |
| GPT-4o [openai2024gpt4o] | |||
| Qwen3-VL-Plus [qwen2025vl] | |||
| Doubao-Seed-1.6-Vision [doubao2024seed] | |||
| Qwen3-VL-4B-Instruct [qwen2025vl_4b] |
3.2 Human-centric World Model (HWM)
| CogVideo [hong2022cogvideo] | DynamiCrafter [xing2024dynamicrafter] | |||
| Metric | w/o WiYH | w/ WiYH | w/o WiYH | w/ WiYH |
| Consistency | 88.2 (+7.6) | 91.5 (+6.2) | ||
| Smoothness | 98.4 (+4.9) | 98.6 (+1.2) | ||
| Dynamic | 78.5 (+15.6) | 89.6 (+14.8) | ||
| Quality | 67.6 (+7.5) | 70.6 (+7.1) | ||
3.2.1 Language-conditioned Video Generation
This section evaluates the contribution of our proposed WiYH dataset to enhancing video generation performance in human-centric manipulation tasks. We fine-tuned two baseline models, CogVideo [hong2022cogvideo] and DynamiCrafter [xing2024dynamicrafter], initialized with their original weights, using WiYH under the image-and-text-to-video setting. During fine-tuning, all input videos were resized to 480p resolution.
As shown in Figure 5, models fine-tuned with WiYH generate egocentric videos with improved spatial and temporal coherence, as well as stronger alignment with both the input image and action instructions. For example, in the “Wipe the cabinet” task, the cabinet’s orientation and the towel’s color are reproduced more accurately. In contrast, videos generated by non-fine-tuned models exhibit noticeable object fragmentation and distortion, along with discontinuous and physically implausible hand–object interactions. Notably, after fine-tuning, the complex motions of non-rigid objects in the generated videos appear more realistic and natural. These improvements demonstrate that WiYH substantially enhances a model’s ability to capture real-world dynamics.
Quantitative results are summarized in Table 3. We use four metrics from VBench [huang2024vbench]—consistency, smoothness, dynamic (motion intensity), and quality—to assess motion plausibility, temporal coherence, and visual fidelity. After fine-tuning with the WiYH dataset, both baseline models show clear improvements across all metrics, with the most significant gains observed in motion intensity and temporal consistency.
3.2.2 Gaussian Splatting 4D Reconstruction
This subsection evaluates the dataset’s utility for supervised end-to-end 4D reconstruction using Gaussian Splatting (GS). By providing high-quality RGB images, depth maps, and camera poses, WiYH supports shape and motion modeling for dynamic scenes. We extracted geometry using MegaSAM [megasam], tracked pixel trajectories of dynamic objects with TAPIR [tapir], and performed scene reconstruction with Shape of Motion [shapeofmotion]. The successful application of our dataset in spatial perception tasks like 4D Gaussian Splatting (4DGS) demonstrates the critical role of its rich annotations, particularly depth information, in enhancing spatial understanding and representation.
As shown in Table 4, photometric metrics remain relatively consistent, whereas geometric metrics vary notably across tasks. These differences highlight the inherent challenges in reconstructing dynamic scenes, where accurate geometry prediction is particularly complex in motion-rich environments. This observation suggests promising directions for future research in 4D reconstruction for robotic manipulation, particularly methods that improve geometric understanding in complex dynamic scenes.
Figure 6 presents the 4D reconstruction results at various time steps, showcasing the dataset’s potential for dynamic scene modeling. By capturing high-quality dynamic scenes, WiYH can help leverage Gaussian Splatting (GS) as an effective 3D tokenizer, advance spatial understanding, and enrich Real2Sim pipelines.
| Scene | Task | Photometry | Geometry | |||
| PSNR | SSIM | LPIPS | A.R. | |||
| 1 | Pouring Wine | 20.15 | 0.815 | 0.206 | 1.04 | 52.01 |
| 2 | Packing Clothes | 20.02 | 0.656 | 0.423 | 2.89 | 17.83 |
| 3 | Packing Cargo | 18.83 | 0.749 | 0.552 | 3.52 | 13.79 |
3.3 Human-centric VLA Training (HVT)
In this section, we evaluate several Vision-Language-Action (VLA) algorithms trained on the WiYH dataset to explore whether its visual-action-language tri-modal support enables effective training of vision-language-action-aligned manipulation policies.
Specifically, we use GR00T [bjorck2025gr00t] and OpenVLA-OFT [kim2025fine] as baseline algorithms, training each to predict the future 1.6-second motion trajectory (16 points at 10 Hz) of the human wrist from the dataset. We sample 10% of the data from each scenario for training and assemble a test set of 400 non-overlapping episodes. Each baseline model is trained for 30,000 steps with a total batch size of 128 on 8 NVIDIA A100 GPUs.
As shown in Table 5, training with human-centric data consistently reduces the error in tracking the ground-truth trajectories on the test set across both mainstream VLA frameworks. These results demonstrate WiYH’s effectiveness for training VLA algorithms. We believe the WiYH dataset can serve as a foundational resource for validating scaling laws in human-centric VLA pre-training.
| Model | Mean Error (cm) | ||
| 0 Steps | 15k Steps | 30k Steps | |
| OpenVLA-OFT [kim2025fine] | 4.75 | 1.28 | 1.12 |
| GR00T [bjorck2025gr00t] | 4.54 | 1.31 | 1.07 |
4 Cross-Embodiment Analysis
In this section, we conduct co-training experiments using cross-embodiment data to investigate the effectiveness of the human video data.
4.1 Implementation Details
We set up two scenarios: a single-object scene and a multi-object cluttered scene. In this experiment, we employ a co-training policy using data collected by an improved version of DexUMI [xu2025dexumi] and from Human Video demonstrations. Figure 8 shows representative scene examples and the data collection setup. For UMI, data collection always started from a limited number of initial poses. The improved UMI system is capable of capturing trajectories with millimeter-level accuracy and was exclusively used to collect high-precision manipulation data in the single-object scene, resulting in fine-grained motion trajectories. In contrast, Human Video contains only manipulation sequences from multi-object cluttered scenes, providing data that spans a larger action space, exhibits a broader distribution of motions, and covers a greater diversity of object observations.
4.2 Results and In-depth Discussion
In the single-object scenario, the robot arm begins execution from a random initial pose during testing, which differs from the initial pose used in the UMI dataset. In the multi-object cluttered scenario, we primarily evaluate the performance of the co-training strategy under conditions of varying scene height, changes in the target object’s pose, and increased occlusion. The configurations of these test scenarios and qualitative experimental results are illustrated in Figure 7.
Quantitative results are summarized in Table 6. (1) In the single-object scenario, co-training the policy with both UMI data and Human Video data leads to an improvement in the success rate of more than 13%. Although the motion precision in the human data is limited, its diverse action distribution enhances the policy’s generalization capability to unseen initial states. (2) Policies trained solely on UMI data exhibit almost no generalization ability in complex scenarios; however, after co-training with human data, the SR increases by 52%. This demonstrates that incorporating human data significantly improves the policy’s ability to interpret complex scenes. Therefore, the co-training approach effectively introduces both the action domain and observation domain of human data into the robot dataset, thereby enhancing the generalization capability of the learned policy.
| Scenario | Data Sources and Amounts | SR (%) | |
| UMI (clips) | Human Video (clips) | ||
| Single-object Scene | 200 | 0 | 20 |
| 200 | 800 | 23.3 (+3.3) | |
| 500 | 0 | 43.3 | |
| 500 | 800 | 56.7 (+13.4) | |
| Cluttered Scene | 200 | 0 | 0.0 |
| 200 | 800 | 32.0 (+32.0) | |
| 500 | 0 | 8.0 | |
| 500 | 800 | 60.0 (+52.0) | |
5 Conclusion
This paper presents the WiYH ecosystem, a cornerstone for embodied intelligence research in real-world environments. We make three key contributions: the Oracle Suite, a wearable system enabling in-the-wild data capture with integrated markerless auto-labeling; the large-scale WiYH Dataset, offering diverse, skilled human manipulation sequences with rich multi-modal streams; and a suite of Comprehensive Benchmarks, supported by extensive atomic-action and vision-language annotations, for holistic evaluation from perception to physical interaction. Our benchmarks reveal that current models still lack the fine-grained spatial and causal reasoning required for robust embodied operation. The WiYH dataset, with its unique scale, and rich annotations, is designed to bridge this gap. We believe this work provides an essential foundation for developing more capable and generalizable embodied agents.
Appendix A World in Your Hand (WiYH) Dataset
A.1 Detailed Task-Skill Counting Statistics
We further supplement the visualization of the number of skills corresponding to each task. Due to space limitations, we report statistics based on only 25% of the dataset, as shown in Figure. 16. The full visualization analysis will be updated subsequently on the dataset website.
A.2 Additional annotation examples visualizations
We provide additional annotation visualizations for the remaining 6 scenes, as shown in Figure 15.
A.2.1 Data Annotation Details.
Raw videos are structured into a three-tier hierarchy: Scene - Task - Subtask by Annotators. Each subtask corresponds to the minimal object-centered action unit or a temporally cohesive action sequence. (1) Subtask Annotation:
Actions: Drawn from predefined operation sets (e.g., PICK, PLACE_INTO, SEARCH).
Target Object: Semantic structure describing action objects and positional relationships (e.g., [object, container]).
Temporal Bounds: Precise start/end timestamps delimiting the execution segment in the video.
Status: Labeled as success, failure, irrelevant to denote execution outcome.
(2) Subtask Prediction chain-of-thought (CoT): With the subtask annotation, we select tasks of moderate length (containing 4-10 subtasks). We annotate them with Chain-of-Thought data, providing reasoning that decomposes the task into several logically connected subtasks. The annotation pipeline for CoT is illustrated in Figure 10. Initially, we provide the model (Qwen2.5-VL) with both the current subtask information (video and text) and the ground-truth text of the next subtask. The model is then prompted to perform posterior attribution, explaining why the specified next subtask logically follows. Subsequently, the attribution is appended to the input and presented again to Qwen2.5-VL, which synthesizes it into a structured causal reasoning process. Finally, the CoT segment is extracted and used to prompt Qwen-3, which is instructed to predict an answer. This predicted answer is matched against the ground-truth answer to validate the completeness and soundness of the CoT labeling.
Appendix B Benchmark Details
B.1 Human-centric Vision-Language
B.1.1 Benchmark Task Descriptions.
The Human-centric Vision-Language Benchmark in Section 3.1 is constructed by sampling raw data from WiYH, followed by manual annotation, including Spatial Referring, Subtask Prediction, and Completion Verification, with distributions illustrated in Figure 9. (1) The Spatial Referring is designed as a spatial region grounding form: given relevant image frames, annotators delineate ROI areas using polygons according to the given questions. It evaluates spatial understanding by prompting the model to output a set of points representing the spatial region referenced in the question. The primary metric is the proportion of these points that fall within the ground-truth polygon. (2) The Subtask Prediction adopts a multiple-choice question format. Using Qwen2.5-VL-72B, we first generate 9 negative candidate options based on video segments and ground truth, spanning action, object, and orientation dimensions. Annotators then manually select the top three hard negative samples, which, together with the correct answer, form the final candidate set. (3) The Completion Verification is designed as binary questions: given a video segment randomly truncated from the end, annotators determine whether the subtask has been completed. And this evaluation metric is the selection accuracy rate.
B.1.2 Benchmark Tasks Visualization.
Examples of the three tasks are shown in Figure 11.
B.2 Human-centric World Model
B.2.1 Language-conditioned Video Generation
Baseline. We validate the effectiveness of our WiYH dataset using video generation baselines with different architectures, including transformer (CogVideo [hong2022cogvideo]) and U-Net (DynamiCrafter [xing2024dynamicrafter]).
Evaluation Metrics. Following VBench [huang2024vbench], the evaluation metrics are as follows:
Consistency. To evaluate the temporal consistency of the target manipulation subjects, we calculate the DINO [caron2021emerging] feature similarity across frames. For assessing the temporal consistency of the background scenes, we compute the CLIP [radford2021learning] feature similarity across frames. The overall consistency score is defined as the arithmetic mean of the subject consistency and the background consistency.
Smoothness. To assess the temporal continuity of the generated videos, we employ a motion-consistency metric based on the AMT [li2023amt] video frame interpolation model. Given a generated video with frames , we first remove all odd-indexed frames to obtain . The AMT model is then used to interpolate the missing frames, producing . Finally, we compute the mean absolute error between the interpolated and original frames, where a larger value indicates smoother and more temporally coherent motion.
Dynamic. To prevent the model from producing static videos in pursuit of higher temporal consistency scores, we employ RAFT [teed2020raft] to estimate optical flow strengths between consecutive frames of each generated video. The average of the largest 5% optical flows is used to determine whether the video is static. The final dynamic score is calculated by measuring the proportion of non-static videos generated by the model.
Quality. To assess the visual quality of generated video frames, we employ the MUSIQ [ke2021musiq] image quality predictor trained on the SPAQ [fang2020perceptual] dataset. This metric quantifies common distortions such as overexposure, noise, and blur, with a higher score indicating fewer artifacts.
Implementation Details. During fine-tuning, all input videos are resized to 480p resolution. For CogVideo, we use the pre-trained image-to-video version with a temporal length of 49, a learning rate of 2e-5, and a batch size of 16. For Dynamicrafter, we leverage the model pre-trained at 512 resolution, setting the temporal length to 48, the learning rate to 1e-5, and the batch size to 16.
B.2.2 Gaussian Splatting 4D Reconstruction
Evaluation Metrics. In terms of metrics of 4D reconstruction, we evaluate the RGB rendering quality using PSNR, SSIM, and LPIPS. For PSNR and SSIM, higher is better, while for LPIPS, lower is better. For geometry accuracy from rendered depth maps, we use A.R. (lower is better) and (higher is better).
PSNR (Peak Signal-to-Noise Ratio),
where is the reference image, is the rendered image, and is the maximum possible pixel value. A higher PSNR indicates that the rendered image is closer to the ground truth.
SSIM (Structural Similarity Index Measure).
where , , and denote the mean, variance, and covariance of image patches and , respectively. SSIM evaluates perceptual similarity in terms of luminance, contrast, and structure; values closer to indicate better similarity.
LPIPS (Learned Perceptual Image Patch Similarity). This metric computes the perceptual distance between rendered and ground-truth images by extracting deep features from a pretrained CNN and measuring the distance in that feature space. Lower LPIPS values indicate that two images are more perceptually similar.
AbsRel (Absolute Relative Error).
where is the predicted depth at the pixel and is the corresponding ground-truth depth. A lower value denotes a more accurate geometry.
(Threshold Accuracy / Delta1).
which measures the percentage of predicted depth values that are within a factor of of the ground-truth depth. A higher value indicates better accuracy.
Implementation Details. During optimization, all input videos are processed at a resolution of 960×720. We represent each scene with 40k dynamic Gaussians and 100k static Gaussians and use B = 10 shared SE(3) motion bases for all points. We optimize the model with Adam, running 1k iterations for the initialization stage and 500 training epochs, with learning rates of for Gaussian means and SE(3) bases, for scales, for rotations, and for colors, opacities, and motion coefficients. Training on a 300-frame sequence takes about 2 hours on a single A100 GPU and supports real-time rendering at around 140 FPS.
B.3 Details of Cross-Embodiment Analysis.
In this section, we present the details of our cross-embodiment experiments in Section 4, Cross-Embodiment Analysis, of the main text. As shown in Figure 14, we demonstrate 11 task variants under two scenarios: Single-Object Scene and Multi-Object Cluttered Scene. Figures 12 and 13 illustrate the success rates and failure causes for each task in these two scenarios, respectively.
As shown in Figures 12 and 13, on most tasks, the incorporation of human data leads to a general increase in success rate. In the Single-Object Scene, the diverse action distribution of human data enhanced the policy’s generalization capability to unseen initial states, enabling successful grasping at different positions or over long distances. In the Multi-Object Cluttered Scene, policies trained solely on UMI data demonstrated almost no generalization ability in complex scenarios. When factors such as object height or layout changed, policies trained only on UMI data failed, whereas the inclusion of human data gradually enabled success by effectively introducing both the action and observation domains of human data into the robot dataset.
Furthermore, the consistently low success rates across all methods in certain tasks highlight important directions for future research: how to further improve the collection precision of human data and how to design efficient algorithms that can better transfer knowledge from human data to dexterous manipulation tasks.
Appendix C Ethics and Privacy Statement
Our work involves large-scale data collection of real-world human manipulation activities. We took a number of measures to ensure responsible data handling, privacy protection, and ethical compliance.
Data Collection and Consent. All human subjects participating in data collection were recruited voluntarily and provided informed consent in accordance with institutional guidelines. Participants were briefed on the purpose of the study, the modalities collected (RGB video, depth, motion capture, tactile sensing), the intended research use, and their right to withdraw at any time.
Privacy Protection. Although the dataset contains real-world human activities and may capture visible human faces or identifiable biometric patterns, several privacy-preserving strategies were implemented:
Controlled subjects only: No bystanders or non-consenting individuals appear in the dataset; all scenes were recorded in controlled environments with only consenting operators present.
No identity labels: We do not include personal information, demographic information, or identity annotations.
Face and identity protection: When faces are visible in RGB streams, we apply optional face blurring in the public release version.
No audio recordings: We do not collect speech or audio signals to avoid unintended disclosure of personal information.
Sensitive Content and Safety. The dataset does not include minors, hazardous behaviors, or sensitive activities. All recordings were conducted in standard workplace or daily-life environments using safety protocols.
Data Usage and Licensing. The released dataset is intended exclusively for research on embodied AI, manipulation, perception, and VLA learning. Redistribution or commercial use is restricted according to the license accompanying the dataset. Users are required to comply with ethical research practices and avoid any attempts at identity recognition or misuse of human-related data.
Potential Risks and Mitigations. As with any dataset containing human recordings, there is a theoretical risk of model misuse, such as identity extraction or surveillance applications. To mitigate these risks, we provide: (1) no textual identity metadata, (2) optional visual anonymization tools, and (3) a license that explicitly forbids re-identification, face recognition, or surveillance-related uses.
Institutional Review. All data collection procedures were reviewed and approved by an internal ethics committee to ensure compliance with privacy protection and responsible AI principles.

### v3

@@
 Abstract
-Large-scale pre-training is fundamental for generalization in language and vision models, but data for dexterous hand manipulation remains limited in scale and diversity, hindering policy generalization.
-Limited scenario diversity, misaligned modalities, and insufficient benchmarking constrain current human manipulation datasets.
-To address these gaps, we introduce World In Your Hands (WiYH), a large-scale open-source ecosystem for human-centric manipulation learning.
-WiYH includes (1) the Oracle Suite, a wearable data collection kit with an auto-labeling pipeline for accurate motion capture; (2) the WiYH Dataset, featuring over 1,000 hours of multi-modal manipulation data across hundreds of skills in diverse real-world scenarios; and (3) extensive annotations and benchmarks supporting tasks from perception to action.
-Furthermore, experiments based on the WiYH ecosystem show that integrating WiYH’s human-centric data significantly enhances the generalization and robustness of dexterous hand policies in tabletop manipulation tasks.
-We believe that World In Your Hands will bring new insights into human-centric data collection and policy learning to the community.
+We introduce World In Your Hands (WIYH), a large-scale open-source ecosystem comprising over 1,000 hours of human manipulation data collected in-the-wild with millimeter-scale motion accuracy.
+Specifically, WIYH includes (1) the Oracle Suite, a wearable data collection kit with an auto-labeling pipeline for accurate motion capture; (2) the WIYH Dataset, featuring over 1,000 hours of multimodal manipulation data across hundreds of skills in diverse real-world scenarios; and (3) extensive annotations and benchmarks supporting tasks from perception to action.
+Furthermore, experiments based on the WIYH ecosystem show that integrating WIYH’s human-centric data improves robotic manipulation success rates from 8% to 60% in cluttered scenes.
+World In Your Hands provides a foundation for advancing human-centric data collection and cross-embodiment policy learning.
+All data and hardware design will be open-source.
 1 Introduction
-| Dataset | Hours | Clips | Wild | Skills | Cam. Calib | RGB | Tactile | Action | Depth | Mask | Instruct. | Reason |
-| Ego4D [grauman2022ego4d] | 3,670 | - | ✓ | ✓ | ✗ | ✓ | ✗ | 2D | ✗ | ✗ | ✓ | ✓ |
-| HOI4D [liu2022hoi4d] | 44.4 | 4k | ✓ | ✓ | ✓ | ✓ | ✗ | 2D | ✓ | ✓ | ✗ | ✗ |
-| HoloAssist [wang2023holoassist] | 166 | 2.22k | ✗ | ✗ | ✗ | ✓ | ✗ | - | ✓ | ✗ | ✓ | ✓ |
-| Ego-Exo4D [grauman2024ego] | 1,286 | 5.04k | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
-| CaptainCook4D [peddi2024captaincook4d] | 94.5 | 384 | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ |
-| HOT3D [banerjee2025hot3d] | 13.9 | 425 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✗ | ✗ | ✗ | ✗ |
-| HO-Cap [wang2024ho] | - | 64 | ✗ | ✗ | ✓ | ✓ | ✗ | 3D | ✓ | ✓ | ✗ | ✗ |
-| Ego-ViD-5M [wang2024egovid] | - | 5M | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
-| EgoDex [hoque2025egodex] | 829 | 338k | ✗ | ✓ | ✗ | ✓ | ✗ | 3D | ✗ | ✓ | ✓ | ✗ |
-| WiYH (Ours) | 1045 | 125.4k | ✓ | ✓ | ✓ | ✓ | ✓ | 3D | ✓ | ✓ | ✓ | ✓ |
-Large-scale pre-training has been established as a cornerstone for achieving generalization in Large Language Models [brown2020language, chowdhery2023palm, achiam2023gpt, touvron2023llama, yang2025qwen3], Vision-Language Models [radford2021learning, li2022blip, alayrac2022flamingo, li2023blip, li2024llava] and Vision-Language-Action Models [brohan2022rt, zitkovich2023rt, kim2024openvla, black2410pi0, wen2024diffusion, zhou2025vision].
-However, datasets for dexterous hand manipulation remain far smaller and less diverse than language datasets, limiting progress in learning robust manipulation policies.
+Large-scale pre-training has been established as a cornerstone for achieving generalization in Large Language Models [brown2020language, chowdhery2023palm, achiam2023gpt, touvron2023llama, yang2025qwen3], Vision-Language Models [radford2021learning, li2022blip, alayrac2022flamingo, li2023blip, li2024llava] and Vision-Language-Action (VLA) Models [brohan2022rt, zitkovich2023rt, kim2024openvla, black2410pi0, wen2024diffusion, zhou2025vision].
+However, datasets for dexterous hand manipulation remain far smaller and less diverse than language datasets, limiting progress in learning robust manipulation.
 To address the problem of data scarcity, human-centric data collection and learning have attracted growing interest [hsieh2025dexman, zhu2025learning, routray2025vipra, zhu2025let, kareer2025egomimic].
 Some approaches [chao2021dexycb, kwon2021h2o, banerjee2025hot3d] capture human hand data during object manipulation using devices such as VR systems and smart glasses, while others [grauman2022ego4d, xu2021videoclip, wang2021actionclip, he2024egovm] leverage foundation models to extract human actions from unlabeled videos of human manipulation available on the internet.
+| Dataset | Hours | Clips | Wild | Cam. Calib | RGB | Tactile | Action | Depth | Mask | Instruct. | VLM |
+| Ego4D [grauman2022ego4d] | 3,670 | ✓ | ✓ | ✗ | ✓ | ✗ | 2D | ✗ | ✗ | ✓ | ✓ |
+| HOI4D [liu2022hoi4d] | 44.4 | 4k | ✓ | ✓ | ✓ | ✗ | 2D | ✓ | ✓ | ✗ | ✗ |
+| HoloAssist [wang2023holoassist] | 166 | 2.22k | ✗ | ✗ | ✓ | ✗ | - | ✓ | ✗ | ✓ | ✓ |
+| Ego-Exo4D [grauman2024ego] | 1,286 | 5.04k | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ |
+| CaptainCook4D [peddi2024captaincook4d] | 94.5 | 384 | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ |
+| HOT3D [banerjee2025hot3d] | 13.9 | 425 | ✗ | ✓ | ✓ | ✗ | 3D | ✗ | ✗ | ✗ | ✗ |
+| HO-Cap [wang2024ho] | - | 64 | ✗ | ✓ | ✓ | ✗ | 3D | ✓ | ✓ | ✗ | ✗ |
+| Ego-ViD-5M [wang2024egovid] | - | 5M | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
+| EgoDex [hoque2025egodex] | 829 | 338k | ✓ | ✗ | ✓ | ✗ | 3D | ✗ | ✓ | ✓ | ✗ |
+| WIYH (Ours) | 1045 | 125.4k | ✓ | ✓ | ✓ | ✓ | 3D | ✓ | ✓ | ✓ | ✓ |
 Despite these advances, current human manipulation datasets still face three major limitations, as shown in Table 1 that hinder their usefulness for dexterous manipulation research: (1) Limited Scenario Diversity: Many datasets are collected in constrained laboratory environments, lacking the diversity and complexity of real-world settings.
-(2) Inadequate Alignment: Existing collections often lack one or more key modalities, such as the absence of fine-grained language instructions, precise 3D hand poses, or object poses, resulting in misalignment across vision, language, and action.
+(2) Inadequate Alignment: Existing collections often lack one or more key modalities, such as the absence of fine-grained language instructions, precise 3D action, or object poses, resulting in misalignment across vision, language, and action.
 (3) Insufficient Benchmarking: Most datasets do not support comprehensive evaluation benchmarks that span the entire pipeline of dexterous manipulation, including scene perception, task planning, and action.
@@
 Specifically, the WIYH ecosystem comprises three core components:
-The Oracle Suite and Auto-labeling Pipeline.
-We introduce a lightweight and flexible wearable data collection kit, the Oracle Suite, designed to capture data without impeding the operator’s natural movements.
-An integrated auto-labeling pipeline enables markerless self-localization and centimeter-level motion accuracy without requiring external third-person tracking.
-The WiYH Dataset: A Large-Scale Multi-modal Human-centric Dataset.
-We present the WiYH dataset, a large-scale multi-modal resource comprising over 1,000 hours of manipulation sequences across hundreds of skills, captured in 10 diverse industrial and daily-life scenarios such as factories, supermarkets, hotels, and restaurants.
-The dataset includes rich sensory streams: multi-view images, camera calibration parameters, depth maps, 3D hand poses, and 6D wrist trajectories.
-Diverse Annotations and Comprehensive Benchmarks.
-Building upon the WiYH dataset, we provide extensive annotations, including approximately 400 hours of atomic-level action instructions and 100k instances of vision-language data for task reasoning, affordance understanding, and object localization.
-These annotations support a comprehensive set of benchmarks—spanning visual language understanding, world model evaluation, and vision-language-action tasks—designed to assess capabilities from scene perception to embodied operation holistically.
-Based on the WiYH ecosystem, we study how human-centric data empowers dexterous manipulation learning in tabletop manipulation tasks.
-We train models on mixtures of Oracle Suite demonstrations and other manipulation data and evaluate the task success rate across varied environments and object layouts.
-Our experiments demonstrate that human-centric data substantially improves the generalization and robustness of dexterous hand policies.
-[Limitation].
-All manipulation tasks are evaluated in real-world environments.
-To increase accessibility and reproducibility, we are developing high-fidelity simulation environments and plan to release them in a future version of the dataset and benchmarks.
+The Oracle Suite.
+We introduce a lightweight, flexible wearable data-collection kit that captures 3D action data with millimeter-scale translational accuracy (mean error below 5 mm) without external third-person tracking.
+The WIYH Dataset.
+We present a large-scale multi-modal resource comprising over 1,000 hours of manipulation demonstrations across 100 skills, captured in 10 diverse scenarios.
+The dataset includes rich sensory streams: multi-view images, camera calibration parameters, hand poses, and wrist trajectories.
+Diverse Annotations.
+Building upon the WIYH dataset, we provide extensive annotations, including 400 hours of atomic action instructions and 100k instances of vision-language data, supporting a comprehensive usage from scene perception to robotic manipulation.
+Based on the WIYH ecosystem, we study how human-centric data empowers dexterous manipulation learning in robotic manipulation tasks.
+We conduct two sets of experiments: (1) cross-embodiment pretraining and (2) retargeting co-training, and evaluate the task success rate across varied environments and object layouts.
+Our experiments demonstrate that human-centric data substantially improves the generalization and robustness of robotic manipulation, whether as large-scale cross-embodiment pre-training data or retargeted action data.
+Limitation. To increase accessibility and reproducibility, we are developing high-fidelity simulation environments and plan to release them in a future version of the dataset and benchmarks.
 2 World In Your Hands Ecosystem
 2.1 Oracle Suite: Human-centric Data Collection Suite
-To enable large-scale, high-fidelity, human-centric multimodal data acquisition, we developed Oracle Suite, a low-cost wearable data collection system designed for human-centric embodied AI research.
-Specifically, as shown in Figure 2, Oracle Suite comprises a hardware system for acquiring multi-view RGB, IMU-based localization, and tactile information, alongside an autolabeling module that achieves high-precision motion capture through the fusion of IR, RGB, and IMU localization.
+We develop Oracle Suite, a low-cost wearable data collection system designed for human-centric multimodal data collection.
+Specifically, as illustrated in Figure 2, Oracle Suite comprises a hardware system that captures multi-view RGB images, IMU-based localization, and tactile information and an autolabeling module that achieves high-precision motion capture by fusing IR, RGB, and IMU localization.
 Hardware System Overview.
-Oracle Suite adopts a modular architecture integrating three primary hardware components: H-HPVHive, H-Gloves, and H-Backpack.
-This modular design enables flexible deployment across diverse environments, without being restricted to controlled indoor labs.
-The Oracle Suite introduces major optimizations in ergonomics, sensing, and reliability:
-H-FPVHive: It supports chest-mounted configuration, and integrates two fisheye cameras, two pinhole cameras, and four infrared lenses.
-All cameras are hardware synchronized.
-The infrared lenses are used for better locating the relative poses of H-Gloves.
-H-Gloves: Each glove contains 6 IMUs, 5 pressure sensors, and an on-board MCU.
-The gloves can capture the hand motions and tactile.
-Moreover, it also contains 3 fisheye cameras for visual tracking and observation.
-H-Backpack: It provides data storage, computation, and power supply. The computation is hosted by NVIDIA Orin.
+Oracle Suite adopts a modular architecture integrating three primary hardware components: H-FPVHive, H-Gloves, and H-Backpack.
+This design enables flexible deployment across diverse environments without being restricted to stable indoor labs.
+Hardware Design. The Oracle Suite incorporates significant optimizations in ergonomics, sensing, and reliability.
+H-FPVHive supports a chest-mounted configuration with two fisheye cameras, two pinhole cameras, and four infrared lenses integrated into the chest unit, where all cameras achieve hardware synchronization, and the infrared lenses facilitate better localization of the relative pose of H-Gloves.
+H-Gloves include six IMUs, five fingertip pressure sensors with a resolution of 5mN and a range of 0.2 N to 50 N, and an onboard MCU per glove to capture hand motions and tactile information while also containing three fisheye cameras for visual tracking and observation.
+H-Backpack provides data storage, computation, and power supply, utilizing an NVIDIA Orin computing core.
 Autolabeling Pipeline.
-As shown in the right part of Figure 2, the Autolabeling Pipeline comprises both online and offline algorithms.
-It fuses multi-modal sensor data (IR + IMU + RGB) to ensure positioning accuracy in in-the-wild scenarios.
-Specifically, the online localization module performs lightweight coarse positioning using visual-inertial odometry (VIO) algorithms and IR detection, while the offline stage refines these coarse estimates through visual Structure-from-Motion (SfM) algorithms.
-The pipeline ultimately outputs centimeter-level 6D poses for both wrists.
-Data Production Efficiency.
-The Oracle Suite pipeline follows a collection–annotation–inspection–archiving workflow for scalable data production.
-A single collector operating 5 hours per day generates approximately 1.8 TB of valid multimodal data.
-At scale, 100 collectors over 30 days yield roughly 5.4 PB of raw data (stored in cloud repositories).
-Efficiency is further improved through standardized capture protocols and automated cloud-based annotation.
-2.2 Data Annotation Pipeline
-Data collection: Practitioners wore the Oracle Suite in each scene to record natural manipulation behavior.
-Each recording session lasted approximately 10 minutes and included start and end event markers.
-The collected signals comprise tactile signals from pressure sensors, binocular fisheye images from the left and right wrist mounts, first-person binocular fisheye images, left-hand and right-hand IMU localization data, and the corresponding camera calibration parameters.
+The autolabeling pipeline involves online and offline algorithms, as shown on the right side of Figure 2.
+By fusing multimodal sensor streams (IR, IMU, and RGB), it maintains robust localization in in-the-wild scenarios and ultimately outputs 6D wrist pose trajectories with millimeter-scale translational accuracy (mean positional error below 5 mm).
+Efficiency of Oracle Suite.
+Compared to teleoperation, Oracle Suite achieves 5 higher collection efficiency (720 vs.
+150 episodes/day) without extensive operator training.
+Compared to VR-based hand tracking, Oracle Suite provides more accurate 3D hand skeletons, particularly under visual occlusion.
+2.2 Data Validation System
+We employ motion capture tests and projection intersection methods to validate data quality.
+Figure 3 illustrates the testing effects of both methods, and we retain valid high-quality data.
+For the motion capture test, we attach motion capture spheres to H-Gloves and complete operations identical to those in the WIYH dataset within a motion-capture room.
+This setup tests the deviation between the wrist trajectory generated by the Oracle Suite autolabeling pipeline and the motion capture trajectory to verify the effectiveness of the automatic annotation pipeline.
+Our Oracle Suite achieves an average position error of 5 mm in the motion capture room environment.
+For projection overlap, we obtain hand masks in perspective views of H-FPVHive using a hand segmentation algorithm.
+We then project the skeleton data from H-Gloves into the same view and calculate the intersection with the hand mask to determine accuracy.
+3 World In Your Hands Dataset
+3.1 Data Collection and Annotation
+Data collection: During data collection, collectors wear the Oracle Suite in various scenarios to record natural manipulation behaviors.
+Each task is recorded once according to the Standard Operating Procedure of the scenario.
+Recording sessions last approximately 10 minutes and include start and end event markers.
+Captured signals include tactile signals from pressure sensors, wrist images, first-person chest images, corresponding camera calibration parameters, left and right hand trajectory data, and left and right hand skeleton information.
 Samples of data annotations are shown in Figure 6.
-Data annotation: The annotation pipeline comprises three parts: subtask (atomic task) annotation, perception annotation, and vision-language annotation.
-Subtask annotation: To construct a large-scale, high-quality set of atomic tasks, we designed a four-stage hybrid annotation pipeline consisting of data processing, task definition, annotation consolidation, and post-processing:
-(1) Data processing: First-person fisheye videos are segmented into 30-second clips, which serve as the basic units for semantic annotation by the video-language module;
-(2) Task definition (manual annotation): The video-language module parses hierarchical operational semantics from the first-person videos and hierarchically decomposes an operation sequence into high-level tasks and atomic subtask units.
-Each task unit contains a preset action described in natural language and a manually annotated target object.
-(3) Data checking (manual review): Outputs from the video-language module are subject to iterative manual review; any incorrect annotations are reprocessed before proceeding to the consolidation stage.
-(4) Post-processing: Annotated preset-action and target-object names are assembled into complete operation instructions and are further augmented using a large language model (LLM).
-After subtask annotation, the raw data are organized into a hierarchical format of tasks composed of multiple atomic subtasks.
-Perception annotation: To provide richer perceptual information, we pre-annotated scene elements using visual foundation models and then performed manual selection and correction.
-The resulting perception annotations include:
-(1) Segmentation masks: instance masks are generated by applying SAM to human-provided prompt points;
-(2) Depth maps: scene depth is estimated from first-person binocular images using the FoundationStereo model.
-Vision-language annotation: The vision-language annotations were produced through manual curation and comprise the following tasks:
-(1) Spatial Referring: For each subtask, we extract a key frame near the action start that contains the operator’s object of attention and in which the target object is not occluded by the hand, body, or other objects.
-Based on the subtask’s target object, spatial relational terms are added to uniquely refer to a single target region in the image.
-(2) Subtask Prediction: We select tasks that contain 4–5 subtasks.
-The next subtask following the current one is used as the correct option.
-Nine incorrect options are generated by combining scene object information and the task’s preset actions; annotators then select the three incorrect options that are the most logically plausible.
-Additionally, Qwen-2.5VL-72B was used to generate auxiliary “thinking” data for next-action prediction, which were subsequently corrected by human annotators;
-(3) Completion Verification: We select tasks containing 4-5 subtasks and randomly trim the tail of the video by 20%, 25%, 30%, or 35% of frames.
-The task description and the truncated video are provided as input, and human annotators label whether the video depicts a completed task.
-Through the above procedures, we obtain rich annotations covering scene understanding and task planning, which constitute the foundation of a unified benchmark.
+After data collection, we curate the WIYH dataset via a three-stage annotation pipeline (Figure 4), comprising atomic action annotation, perception annotation, and vision-language annotation.
+Atomic action annotation: To construct data suitable for robotic foundation model training, we design a four-stage hybrid annotation process for atomic action annotation that decomposes a task in the standard operating procedure into multiple atomic segments, where each segment is termed a subtask.
+Video Segmentation: The process initiates with video segmentation, where Qwen2.5-VL-72B [Qwen2.5-VL] segments the recorded data for each task into semantic atomic fragments serving as basic units for annotation.
+Task definition: Humans define a preset action termed a "skill" and an operation target object for each atomic fragment.
+Data checking: Manual reviewers check boundary and task definition accuracy to correct errors.
+Instruction Augmentation: Post-processing combines the annotated preset actions and target object names into complete atomic instructions, which are augmented and enhanced using Qwen2.5-VL-72B [Qwen2.5-VL].
+Perception annotation: We use vision foundation models [kirillov2023segany, wen2025foundationstereo] to pre-annotate scene elements, followed by manual selection and correction.
+This includes generating instance masks by applying SAM [kirillov2023segany] to human-provided prompt points and estimating scene depth from first-person binocular images using the FoundationStereo [wen2025foundationstereo] model.
+Vision-language annotation is generated through manual curation and includes several components.
+CoT Reasoning: We use Qwen2.5-VL-72B [Qwen2.5-VL] to generate a thought process indicating the position of each subtask within a complete task and when to enter the next subtask for long-horizon task decomposition, where generated thought data is reviewed and corrected by humans.
+Spatial Referring (SR): We select key frames near the action start where the operation target object is clearly visible for some subtasks and add spatial relation terms based on the atomic instruction to uniquely refer to a single target area in the scene.
+Subtask Prediction (SP): We select tasks containing 4 to 5 subtasks and use the next subtask after the current one as the correct option, while generating three logically reasonable incorrect options based on the task description.
+Completion Verification (CV): We randomly crop 0% to 35% of each subtask for a task and use the atomic instruction of each subtask, as well as the cropped fragment, as input to label the task progress status regarding whether the task is completed.
+Notably, subtask prediction, spatial referring, and progress verification are annotated on only a subset of the whole WIYH data to construct an evaluation set for validating the capabilities of existing VLMs in human-centric scenarios.
 3.2 Data Statistics
@@
 For each scene, we compute the average execution duration of fine-grained actions.
-The dataset exhibits substantial variation across action categories, from short, instantaneous operations such as wipe and insert, to long-horizon manipulation such as smooth out or foldlay.
+The dataset exhibits substantial variation across action categories, from short, instantaneous operations such as "wipe" and "insert" to long-horizon manipulation such as "smooth out" or "fold lay".
 These temporal differences reflect the heterogeneous complexity of real-world household and industrial workflows.
@@
 These distributions further demonstrate that WIYH captures a wide spectrum of real-world manipulation behaviors.
-3 Benchmark and Applications
-3.1 Human-centric Vision-Language (HVL)
+More data, demonstrations, and statistics will be provided in the supplementary materials.
+4 Benchmark and Application
+4.1 Benchmark: Human-centric Vision-Language (HVL)
 In this section, we construct three evaluation tasks based on the vision-language annotations introduced in Section 3.
-These tasks are designed to evaluate whether existing Multimodal Large Language Models (MLLMs) can achieve a genuine understanding of human-centric manipulation tasks in highly generalized in-the-wild environments.
+These tasks are designed to evaluate whether existing Multimodal Large Language Models (MLLMs) can achieve a genuine understanding of human-centric manipulation tasks in highly generalized, in-the-wild environments.
 We define the following three evaluation tasks and test several closed-source and open-source MLLMs, including GPT-4o [hurst2024gpt], Doubao-seed [bai2024seed], and Qwen [yang2025qwen3].
 Detailed experimental setups and data statistics are provided in the supplementary material.
-(1) Spatial Referring (SR): Given a human manipulation image and a description of the spatial relationships among multiple objects in the scene, the model is required to output the pixel coordinates of a specified region.
+(1) Spatial Referring (S.R.): Given a human manipulation image and a description of the spatial relationships among multiple objects in the scene, the model is required to output the pixel coordinates of a specified region.
 Performance is measured by the probability that the predicted point lies within the target area.
-(2) Subtask Prediction (SP): Given a piece of human manipulation video, the current task description, and four candidate options for the next action, the model must select the correct subsequent action.
-(3) Completion Verification (CV): Given a piece of human manipulation video and the current task description, the model needs to determine whether the task has been completed.
-As shown in Table 2, Qwen3-VL-Plus, which was pre-trained on spatial understanding data, achieves the best performance on the Spatial Referring task.
-However, scores of all MLLMs are below 0.5, indicating that current MLLMs still lack sufficient capability in interpreting fine-grained spatial relationships in manipulation scenarios.
-On the Completion Verification task, all models perform slightly above 0.5, suggesting that their ability to accurately judge task completion in such contexts remains limited.
-These experimental results collectively demonstrate that general-purpose, vanilla VLLMs still fall short in embodied tasks.
-Our WiYH dataset, with its rich annotations of human activities and spatial relationships, presents a promising foundation for future embodied pre-training.
-| Model | SR (%) | SP (%) | CV (%) |
+(2) Subtask Prediction (S.P.): Given a piece of human manipulation video, the current task description, and four candidate options for the next action, the model must select the correct subsequent action.
+(3) Completion Verification (C.V.): Given a piece of human manipulation video and the current task description, the model needs to determine whether the task has been completed.
+Evaluation details will be provided in the supplementary material.
+As shown in Table 2, scores for the S.R.
+task and C.V.
+task are generally low, whereas scores for the S.P.
+task remain relatively high.
+This discrepancy reflects the characteristics of existing VLMs, which demonstrate competent performance in high-level task decomposition but exhibit limitations in dynamic progress modeling and spatial understanding in specific embodied tasks.
+Specifically, on the S.R.
+task, scores of all VLMs are below 0.5, and on the C.V.
+task, all models perform slightly above 0.5, suggesting that their abilities in such contexts remain limited.
+Notably, Qwen3-VL-Plus, which was pre-trained on spatial-understanding data, achieves the best performance on the S.R.
+task.
+These experimental results collectively demonstrate that general-purpose, vanilla VLMs still fall short in embodied tasks.
+Our WIYH dataset, with its rich annotations of human activities, scene depth, and object masks, presents a promising foundation for future embodied pre-training.
+| Model | S.R. (%) | S.P. (%) | C.V. (%) |
 | GPT-4o [openai2024gpt4o] | |||
@@
 | Qwen3-VL-4B-Instruct [qwen2025vl_4b] |
-3.2 Human-centric World Model (HWM)
-| CogVideo [hong2022cogvideo] | DynamiCrafter [xing2024dynamicrafter] | |||
-| Metric | w/o WiYH | w/ WiYH | w/o WiYH | w/ WiYH |
-| Consistency | 88.2 (+7.6) | 91.5 (+6.2) | ||
-| Smoothness | 98.4 (+4.9) | 98.6 (+1.2) | ||
-| Dynamic | 78.5 (+15.6) | 89.6 (+14.8) | ||
-| Quality | 67.6 (+7.5) | 70.6 (+7.1) | ||
-3.2.1 Language-conditioned Video Generation
-This section evaluates the contribution of our proposed WiYH dataset to enhancing video generation performance in human-centric manipulation tasks.
-We fine-tuned two baseline models, CogVideo [hong2022cogvideo] and DynamiCrafter [xing2024dynamicrafter], initialized with their original weights, using WiYH under the image-and-text-to-video setting.
-During fine-tuning, all input videos were resized to 480p resolution.
-As shown in Figure 5, models fine-tuned with WiYH generate egocentric videos with improved spatial and temporal coherence, as well as stronger alignment with both the input image and action instructions.
-For example, in the “Wipe the cabinet” task, the cabinet’s orientation and the towel’s color are reproduced more accurately.
-In contrast, videos generated by non-fine-tuned models exhibit noticeable object fragmentation and distortion, along with discontinuous and physically implausible hand–object interactions.
-Notably, after fine-tuning, the complex motions of non-rigid objects in the generated videos appear more realistic and natural.
-These improvements demonstrate that WiYH substantially enhances a model’s ability to capture real-world dynamics.
-Quantitative results are summarized in Table 3.
-We use four metrics from VBench [huang2024vbench]—consistency, smoothness, dynamic (motion intensity), and quality—to assess motion plausibility, temporal coherence, and visual fidelity.
-After fine-tuning with the WiYH dataset, both baseline models show clear improvements across all metrics, with the most significant gains observed in motion intensity and temporal consistency.
-3.2.2 Gaussian Splatting 4D Reconstruction
+4.2 Application: Human-centric World Modeling
 This subsection evaluates the dataset’s utility for supervised end-to-end 4D reconstruction using Gaussian Splatting (GS).
@@
 Figure 7 presents the 4D reconstruction results at various time steps, showcasing the dataset’s potential for dynamic scene modeling.
-By capturing high-quality dynamic scenes, WiYH can help leverage Gaussian Splatting (GS) as an effective 3D tokenizer, advance spatial understanding, and enrich Real2Sim pipelines.
+By capturing high-quality dynamic scenes, WIYH can help leverage Gaussian Splatting (GS) as an effective 3D tokenizer, advance spatial understanding, and enrich Real2Sim pipelines and neural-network-based simulation.
+In addition, we further present the experimental results of video generation in the supplementary materials.
 | Scene | Task | Photometry | Geometry | |||
@@
 | 3 | Packing Cargo | 18.83 | 0.749 | 0.552 | 3.52 | 13.79 |
-3.3 Human-centric VLA Training (HVT)
-In this section, we evaluate several Vision-Language-Action (VLA) algorithms trained on the WiYH dataset to explore whether its visual-action-language tri-modal support enables effective training of vision-language-action-aligned manipulation policies.
-Specifically, we use GR00T [bjorck2025gr00t] and OpenVLA-OFT [kim2025fine] as baseline algorithms, training each to predict the future 1.6-second motion trajectory (16 points at 10 Hz) of the human wrist from the dataset.
-We sample 10% of the data from each scenario for training and assemble a test set of 400 non-overlapping episodes.
-Each baseline model is trained for 30,000 steps with a total batch size of 128 on 8 NVIDIA A100 GPUs.
-As shown in Table 5, training with human-centric data consistently reduces the error in tracking the ground-truth trajectories on the test set across both mainstream VLA frameworks.
-These results demonstrate WiYH’s effectiveness for training VLA algorithms.
-We believe the WiYH dataset can serve as a foundational resource for validating scaling laws in human-centric VLA pre-training.
-| Model | Mean Error (cm) | ||
-| 0 Steps | 15k Steps | 30k Steps | |
-| OpenVLA-OFT [kim2025fine] | 4.75 | 1.28 | 1.12 |
-| GR00T [bjorck2025gr00t] | 4.54 | 1.31 | 1.07 |
-4 Cross-Embodiment Analysis
-In this section, we conduct co-training experiments using cross-embodiment data to investigate the effectiveness of the human video data.
-4.1 Implementation Details
-We set up two scenarios: a single-object scene and a multi-object cluttered scene.
-In this experiment, we employ a co-training policy using data collected by an improved version of DexUMI [xu2025dexumi] and from Human Video demonstrations.
-Figure 8 shows representative scene examples and the data collection setup.
-For UMI, data collection always started from a limited number of initial poses.
-The improved UMI system is capable of capturing trajectories with millimeter-level accuracy and was exclusively used to collect high-precision manipulation data in the single-object scene, resulting in fine-grained motion trajectories.
-In contrast, Human Video contains only manipulation sequences from multi-object cluttered scenes, providing data that spans a larger action space, exhibits a broader distribution of motions, and covers a greater diversity of object observations.
+5 Human-centric Manipulation
+5.1 Experiment Objective and Implementation
+In this section, our objective is to validate the efficacy of the WIYH data collection system and dataset for robotics manipulation.
+To this end, we formulate two core research questions.
+(1) Does utilizing the WIYH dataset as pre-training data enhance the embodied manipulation capabilities of the VLA model?
+(2) Does retargeting WIYH data to dexterous hand actions improve dexterous manipulation performance?
+To address these questions, we design two real-world manipulation experiments: (1) cross-embodiment pretraining and (2) retargeting co-training.
+In the cross-embodiment pretraining experiment, we adopt StarVLA-PI [starvla2025] as the base model.
+We pre-train StarVLA-PI using WIYH data with atomic instruction annotations for both VLM and VLA data.
+For post-training, we select two tasks in a real-world environment: rose insertion and gift packing, and we collect 150 gripper-based teleoperation manipulation samples for each task, as shown on the left side of Figure 8.
+We compare the success rates across three settings, including VLM pre-training, VLA pre-training, and without pre-training.
+VLA pretraining refers to loading all pre-trained VLA weights during post-training, while VLM pretraining denotes loading only the VLM weights.
+During evaluation, each setting of each task is tested 10 times with object positions and target objects varied in each trial.
+In the retargeting co-training experiment, we construct a combined dataset using UMI robot data collected in constrained environments and data of the same task collected using the Oracle Suite.
+The robot system employs a low-degree-of-freedom dexterous hand as the end-effector , and the hand data from is retargeted to match the degrees of freedom of .
+We evaluate the effectiveness of the human-centric data collected by Oracle Suite across various dexterous grasping tasks.
+Implementation details will be provided in the supplementary materials.
+Notably, the right side of figure 8 shows scene examples and the data collection setup.
+The collected robot data always starts from a limited number of initial poses in the single-object scene.
+In contrast, human-centric data contains only manipulation sequences from multi-object cluttered scenes, making a larger observation domain and action space than .
 5.2 Results and In-depth Discussion
-In the single-object scenario, the robot arm begins execution from a random initial pose during testing, which differs from the initial pose used in the UMI dataset.
+In this section, we analyze the results of two experiments to demonstrate how WIYH empowers robotic manipulation.
+For the cross-embodiment pretraining experiments, we demonstrate the quantitative result and the training loss in Table 4 and in Figure 9, respectively.
+As indicated by success rates, utilizing WIYH for VLA pre-training yields the most significant performance improvements, from 15% to 70%.
+Besides, as illustrated in Figure 9, VLA pre-training achieves lower loss and smoother gradient convergence than VLM pre-training and no pre-training.
+This efficacy remains robust given the cross-embodiment gap between the pre-training data (human hands) and post-training data (robot gripper).
+| Method | Rose Insertion | Gift Packing | AVG |
+| w/o Pretrain | 1/10 | 2/10 | 15% |
+| VLM Pretrain | 3/10 | 3/10 | 30% |
+| VLA Pretrain | 6/10 | 8/10 | 70% |
+For the retargeting co-training experiments, we demonstrate the quantitative and qualitative results in Table 5 and in Figure 10, respectively.
+In the single-object scenario, the robot arm begins execution from a random initial pose during testing, which differs from the initial pose used in the robot data .
 In the multi-object cluttered scenario, we primarily evaluate the performance of the co-training strategy under conditions of varying scene height, changes in the target object’s pose, and increased occlusion.
@@
 Quantitative results are summarized in Table 5.
-(1) In the single-object scenario, co-training the policy with both UMI data and Human Video data leads to an improvement in the success rate of more than 13%.
-Although the motion precision in the human data is limited, its diverse action distribution enhances the policy’s generalization capability to unseen initial states.
-(2) Policies trained solely on UMI data exhibit almost no generalization ability in complex scenarios; however, after co-training with human data, the SR increases by 52%.
-This demonstrates that incorporating human data significantly improves the policy’s ability to interpret complex scenes.
-Therefore, the co-training approach effectively introduces both the action domain and observation domain of human data into the robot dataset, thereby enhancing the generalization capability of the learned policy.
-| Scenario | Data Sources and Amounts | SR (%) | |
-| UMI (clips) | Human Video (clips) | ||
+(1) In the single-object scenario, co-training the policy with both robot data and human-centric data leads to an improvement in the success rate of more than 13%.
+Although the motion precision in the human-centric data is limited, its diverse action distribution enhances the policy’s generalization capability to unseen initial states.
+(2) Policies trained solely on robot data exhibit minimal generalization capability in complex scenarios.
+Simply increasing data quantity from 200 clips to 500 clips yields marginal performance improvement from 0% to 8%.
+However, after co-training with human-centric data, the success rate increases by 52%.
+This demonstrates that incorporating human-centric data significantly improves the policy’s ability to interpret complex scenes.
+Therefore, the co-training approach effectively introduces both the action domain and observation domain of human-centric data into the robot dataset, thereby enhancing the generalization capability of the learned policy.
+| Scenario ă | Data Sources and Amounts (Clips) ă | SR (%) | |
+| ă | Robot data amount | Human data amount | ă |
 | Single-object Scene | 200 | 0 | 20 |
@@
 6 Conclusion
-This paper presents the WiYH ecosystem, a cornerstone for embodied intelligence research in real-world environments.
-We make three key contributions: the Oracle Suite, a wearable system enabling in-the-wild data capture with integrated markerless auto-labeling; the large-scale WiYH Dataset, offering diverse, skilled human manipulation sequences with rich multi-modal streams; and a suite of Comprehensive Benchmarks, supported by extensive atomic-action and vision-language annotations, for holistic evaluation from perception to physical interaction.
+In this paper, we present the WIYH ecosystem, a cornerstone for embodied intelligence research in real-world environments.
+We make three key contributions: the Oracle Suite, a wearable system enabling in-the-wild data capture with integrated markerless auto-labeling; the large-scale WIYH dataset, offering diverse, skilled human manipulation sequences with rich multimodal streams; and a suite of comprehensive benchmarks, supported by extensive atomic-action and vision-language annotations, for holistic evaluation from perception to physical interaction.
 Our benchmarks reveal that current models still lack the fine-grained spatial and causal reasoning required for robust embodied operation.
-The WiYH dataset, with its unique scale, and rich annotations, is designed to bridge this gap.
-We believe this work provides an essential foundation for developing more capable and generalizable embodied agents.
-Appendix A World in Your Hand (WiYH) Dataset
-A.1 Detailed Task-Skill Counting Statistics
+The WIYH dataset, with its unique scale and rich annotations, is designed to bridge this gap.
+In the future, we will open-source the whole dataset and hardware design to promote human-centric manipulation research.
+Appendix
+Appendix 0.A World In Your Hands Ecosystem
+0.A.1 Hardware Details
+Camera: We primarily use two fisheye cameras below each wrist and two fisheye cameras on the chest as the raw RGB signal sources.
+These cameras feature a 180° field of view with an original resolution of 1536 × 1920 pixels.
+Tactile: We utilize resistive pressure sensors with a resolution of 5 mN and a range of 0.2 N to 50 N distributed across the fingertips of all fingers.
+Autolabel Pipeline: The autolabeling pipeline involves online and offline algorithms.
+Specifically, the online localization module performs lightweight coarse positioning using visual-inertial odometry (VIO) algorithms [qin2018robust] and IR detection, while the offline stage refines these coarse estimates through Structure-from-Motion (SfM) algorithms [schoenberger2016sfm].
+0.A.2 Data Collection Efficiency
+The Oracle Suite pipeline follows a Collection, Annotation, Inspection, Archiving workflow for scalable data production.
+A single collector operating 8 hours per day generates approximately 1.8 TB of valid multimodal data.
+At scale, 100 collectors over 30 days yield roughly 5.4 PB of raw data (stored in cloud repositories).
+Efficiency is further improved through standardized capture protocols and automated cloud-based annotation.
+We compare the efficiency of different data collection paradigms in Table 6.
+Efficiency: taking the rose insertion task in our experiments as an example, given 8 hours as collection time, teleoperation can collect about 150 data clips; UMI-like methods, such as DexUMI [xu2025dexumi], can collect about 400 data clips; our oracle suite can collect 720 data clips; and VR can collect around 800 data clips.
+Quality: VR-based hand tracking only provides 2D hand skeletons and suffers from visual occlusion.
+In contrast.
+In comparison, the Oracle Suite provides accurate 3D hand skeletons, even if under visual occlusion.
+| Teleop | DexUMI | VR | Oracle Suit (Ours) | |
+| Efficiency | ||||
+| Quality | ||||
+| End-effector | Dexterous | Dexterous | Hands | Hands |
+Appendix 0.B World In Your Hands Dataset
+0.B.1 Detailed Task-Skill Counting Statistics
 We further supplement the visualization of the number of skills corresponding to each task.
-Due to space limitations, we report statistics based on only 25% of the dataset, as shown in Figure.
-16.
+Due to space limitations, we report statistics based on only 25% of the dataset, as shown in Figure 11.
+Please zoom in electronically to inspect the details.
 The full visualization analysis will be updated subsequently on the dataset website.
-A.2 Additional annotation examples visualizations
-We provide additional annotation visualizations for the remaining 6 scenes, as shown in Figure 15.
-A.2.1 Data Annotation Details.
+0.B.2 Data Annotation Details.
 Raw videos are structured into a three-tier hierarchy: Scene - Task - Subtask by annotators.
 Each subtask corresponds to the minimal object-centered action unit or a temporally cohesive action sequence.
-(1) Subtask Annotation:
-Actions: Drawn from predefined operation sets (e.g., PICK, PLACE_INTO, SEARCH).
-Target Object: Semantic structure describing action objects and positional relationships (e.g., [object, container]).
+(1) Atomic (Subtask) Annotation:
+Actions (Skills): Drawn from predefined operation sets (e.g., PICK, PLACE INTO, SEARCH).
+Target Object: Semantic structure describing objects (e.g., [cloth, container]).
 Temporal Bounds: Precise start/end timestamps delimiting the execution segment in the video.
 Status: Labeled as success, failure, irrelevant to denote execution outcome.
-(2) Subtask Prediction chain-of-thought (CoT): With the subtask annotation, we select tasks of moderate length (containing 4-10 subtasks).
+(2) Chain-of-Thought (CoT) Reasoning: With the subtask annotation, we select tasks of moderate length (containing 4-10 subtasks).
 We annotate them with Chain-of-Thought data, providing reasoning that decomposes the task into several logically connected subtasks.
 The annotation pipeline for CoT is illustrated in Figure 12.
-Initially, we provide the model (Qwen2.5-VL) with both the current subtask information (video and text) and the ground-truth text of the next subtask.
+Initially, we provide the model (Qwen2.5-VL-72B [Qwen2.5-VL]) with both the current subtask information (video and subtask instruction) and the ground-truth instruction of the next subtask.
 The model is then prompted to perform posterior attribution, explaining why the specified next subtask logically follows.
-Subsequently, the attribution is appended to the input and presented again to Qwen2.5-VL, which synthesizes it into a structured causal reasoning process.
-Finally, the CoT segment is extracted and used to prompt Qwen-3, which is instructed to predict an answer.
-This predicted answer is matched against the ground-truth answer to validate the completeness and soundness of the CoT labeling.
-Appendix B Benchmark Details
-B.1 Human-centric Vision-Language
-B.1.1 Benchmark Task Descriptions.
-The Human-centric Vision-Language Benchmark in Section 3.1 is constructed by sampling raw data from WiYH, followed by manual annotation, including Spatial Referring, Subtask Prediction, and Completion Verification, with distributions illustrated in Figure 9.
+Subsequently, the attribution is appended to the input and presented again to Qwen2.5-VL-72B, synthesizing it into a structured causal reasoning process.
+Next, the CoT segment is extracted and used to prompt Qwen3 [yang2025qwen3], which is instructed to predict an answer.
+This predicted answer is matched against the ground-truth answer to automatically validate the completeness and soundness of the CoT labeling.
+Finally, human review and correction of erroneous annotations are performed.
+0.B.3 More visualization.
+We present more visualizations about the annotations of our dataset in Figure 13. Annotations of the Human-centric Vision-Language tasks are shown in Figure 14.
+Appendix 0.C Benchmark And Application Details
+0.C.1 Human-centric Vision-Language
+0.C.1.1 Benchmark Task Descriptions.
+The Human-centric Vision-Language Benchmark in Section 4.1 is constructed by sampling raw data from WIYH, followed by manual annotation, including Spatial Referring, Subtask Prediction, and Completion Verification, with the amounts and distributions illustrated in Figure 15.
+In these tasks, all VLMs are directly tested without any training.
 (1) The Spatial Referring is designed as a spatial region grounding form: given relevant image frames, annotators delineate ROI areas using polygons according to the given questions.
-It evaluates spatial understanding by prompting the model to output a set of points representing the spatial region referenced in the question.
-The primary metric is the proportion of these points that fall within the ground-truth polygon.
+Specifically, given the question as prompts, the VLM outputs a set of points representing the spatial region referenced in the question.
+We denote the set of points that fall within the ground-truth region as .
+We evaluate the VLM’s spatial understanding by computing the proportion of the points that fall within the ground-truth region.
+where is the number of and is the number of .
+A is a scalar in the range , with values closer to 1 indicating stronger spatial understanding capability of the VLM.
+This task is designed to evaluate the capability of VLM models to understand spatial referring in human-centric scenarios.
 (2) The Subtask Prediction adopts a multiple-choice question format.
-Using Qwen2.5-VL-72B, we first generate 9 negative candidate options based on video segments and ground truth, spanning action, object, and orientation dimensions.
-Annotators then manually select the top three hard negative samples, which, together with the correct answer, form the final candidate set.
+We use the Qwen2.5-VL-72B tool to first generate nine negative candidate options based on video clips and ground truth values, covering both action and target object dimensions.
+Then, the annotators manually select three reasonable negative samples that fit the current scenario.
+These samples, along with the correct answers, constitute the final candidate set.
+That is, each question has four options, and the accuracy of random selection was approximately 25%.
+This task is designed to evaluate the capability of VLM models to perform long-horizon task planning and decomposition in human-centric scenarios.
 (3) The Completion Verification is designed as binary questions: given a video segment randomly truncated from the end, annotators determine whether the subtask has been completed.
-And this evaluation metric is the selection accuracy rate.
-B.1.2 Benchmark Tasks Visualization.
-Examples of the three tasks are shown in Figure 11.
-B.2 Human-centric World Model
-B.2.1 Language-conditioned Video Generation
-Baseline.
-We validate the effectiveness of our WiYH dataset using video generation baselines with different architectures, including transformer (CogVideo [hong2022cogvideo]) and U-Net (DynamiCrafter [xing2024dynamicrafter]).
-Evaluation Metrics. Following VBench [huang2024vbench], the evaluation metrics are as follows:
-Consistency.
-To evaluate the temporal consistency of the target manipulation subjects, we calculate the DINO [caron2021emerging] feature similarity across frames.
-For assessing the temporal consistency of the background scenes, we compute the CLIP [radford2021learning] feature similarity across frames.
-The overall consistency score is defined as the arithmetic mean of the subject consistency and the background consistency.
-Smoothness.
-To assess the temporal continuity of the generated videos, we employ a motion-consistency metric based on the AMT [li2023amt] video frame interpolation model.
-Given a generated video with frames , we first remove all odd-indexed frames to obtain .
-The AMT model is then used to interpolate the missing frames, producing .
-Finally, we compute the mean absolute error between the interpolated and original frames, where a larger value indicates smoother and more temporally coherent motion.
-Dynamic.
-To prevent the model from producing static videos in pursuit of higher temporal consistency scores, we employ RAFT [teed2020raft] to estimate optical flow strengths between consecutive frames of each generated video.
-The average of the largest 5% optical flows is used to determine whether the video is static.
-The final dynamic score is calculated by measuring the proportion of non-static videos generated by the model.
-Quality.
-To assess the visual quality of generated video frames, we employ the MUSIQ [ke2021musiq] image quality predictor trained on the SPAQ [fang2020perceptual] dataset.
-This metric quantifies common distortions such as overexposure, noise, and blur, with a higher score indicating fewer artifacts.
-Implementation Details.
-During fine-tuning, all input videos are resized to 480p resolution.
-For CogVideo, we use the pre-trained image-to-video version with a temporal length of 49, a learning rate of 2e-5, and a batch size of 16.
-For Dynamicrafter, we leverage the model pre-trained at 512 resolution, setting the temporal length to 48, the learning rate to 1e-5, and the batch size to 16.
-B.2.2 Gaussian Splatting 4D Reconstruction
+The evaluation metric is the selection accuracy rate.
+The accuracy of random selection was approximately 50%.
+This task is designed to evaluate the capability of VLM models to understand the dynamic process of manipulation tasks in human-centric scenarios.
+0.C.1.2 Tasks Visualization.
+We show the samples and detailed annotations of the three tasks in Figure 16.
+0.C.1.3 Case Analysis.
+Figure 16 presents the results of the Qwen3-VL model on the Spatial Referring task.
+We observe that existing VLMs perform reasonably well on in-plane spatial referring relations (e.g., “between A and B” as illustrated in the first row).
+However, when it comes to relations that require understanding positions across different height levels, such as “beneath” or “between A and B” where A and B are situated at different heights, the models fail to grasp the correct spatial information.
+0.C.2 Human-centric World Model
+0.C.2.1 Gaussian Splatting 4D Reconstruction
 Evaluation Metrics.
@@
 which measures the percentage of predicted depth values that are within a factor of of the ground-truth depth. A higher value indicates better accuracy.
-Implementation Details.
 During optimization, all input videos are processed at a resolution of 960×720.
@@
 Training on a 300-frame sequence takes about 2 hours on a single A100 GPU and supports real-time rendering at around 140 FPS.
-B.3 Details of Cross-Embodiment Analysis.
-In this section, we present the details of our cross-embodiment experiments in Section 4, Cross-Embodiment Analysis, of the main text.
+0.C.2.2 Addition Application: Language-conditioned Video Generation
+This section evaluates the contribution of our proposed WIYH dataset to enhancing video generation performance in human-centric manipulation tasks.
+We fine-tuned two baseline models, transformer-based CogVideo [hong2022cogvideo] and UNet-based DynamiCrafter [xing2024dynamicrafter], initialized with their original weights, using WIYH under the image-and-text-to-video setting.
+During fine-tuning, all input videos were resized to 480p resolution.
+We randomly select 90% of the data as the training set and the remaining 10% as the test set.
+As shown in Figure 17, models fine-tuned with WIYH generate egocentric videos with improved spatial and temporal coherence, as well as stronger alignment with both the input image and action instructions.
+For example, in the “Wipe the cabinet” task, the cabinet’s orientation and the towel’s color are reproduced more accurately.
+In contrast, videos generated by non-fine-tuned models exhibit noticeable object fragmentation and distortion, along with discontinuous and physically implausible hand–object interactions.
+Notably, after fine-tuning, the complex motions of non-rigid objects in the generated videos appear more realistic and natural.
+These improvements demonstrate that WIYH substantially enhances a model’s ability to capture real-world dynamics.
+Quantitative results are summarized in Table 7.
+We use four metrics from VBench [huang2024vbench]—consistency, smoothness, dynamic (motion intensity), and quality in order to assess motion plausibility, temporal coherence, and visual fidelity of the video generation models.
+Following VBench [huang2024vbench], the evaluation metrics are as follows:
+Consistency.
+To evaluate the temporal consistency of the target manipulation subjects, we calculate the DINO [caron2021emerging] feature similarity across frames.
+For assessing the temporal consistency of the background scenes, we compute the CLIP [radford2021learning] feature similarity across frames.
+The overall consistency score is defined as the arithmetic mean of the subject consistency and the background consistency.
+Smoothness.
+To assess the temporal continuity of the generated videos, we employ a motion-consistency metric based on the AMT [li2023amt] video frame interpolation model.
+Given a generated video with frames , we first remove all odd-indexed frames to obtain .
+The AMT model is then used to interpolate the missing frames, producing .
+Finally, we compute the mean absolute error between the interpolated and original frames, where a larger value indicates smoother and more temporally coherent motion.
+Dynamic.
+To prevent the model from producing static videos in pursuit of higher temporal consistency scores, we employ RAFT [teed2020raft] to estimate optical flow strengths between consecutive frames of each generated video.
+The average of the largest 5% optical flows is used to determine whether the video is static.
+The final dynamic score is calculated by measuring the proportion of non-static videos generated by the model.
+Quality.
+To assess the visual quality of generated video frames, we employ the MUSIQ [ke2021musiq] image quality predictor trained on the SPAQ [fang2020perceptual] dataset.
+This metric quantifies common distortions such as overexposure, noise, and blur, with a higher score indicating fewer artifacts.
+After fine-tuning with the WIYH dataset, both baseline models show clear improvements across all metrics, with the most significant gains observed in dynamics and temporal consistency.
+| CogVideo | DynamiCrafter | |||
+| Metric | w/o WIYH | w/ WIYH | w/o WIYH | w/ WIYH |
+| Consistency | 88.2 (+7.6) | 91.5 (+6.2) | ||
+| Smoothness | 98.4 (+4.9) | 98.6 (+1.2) | ||
+| Dynamic | 78.5 (+15.6) | 89.6 (+14.8) | ||
+| Quality | 67.6 (+7.5) | 70.6 (+7.1) | ||
+Appendix 0.D Details of Human-centric Manipulation
+In this section, we present the details of our co-training experiments in Section 5, Human-centric Manipulation, of the main text.
 As shown in Figure 18, we demonstrate 11 task variants under two scenarios: Single-Object Scene and Multi-Object Cluttered Scene.
-Figures 12 and 13 illustrate the success rates and failure causes for each task in these two scenarios, respectively.
-As shown in Figures 12 and 13, on most tasks, the incorporation of human data leads to a general increase in success rate.
+Figures 19 and 20 illustrate the success rates and failure causes for each task in these two scenarios, respectively.
+As shown in Figures 19 and 20, on most tasks, the incorporation of human data leads to a general increase in success rate.
 In the Single-Object Scene, the diverse action distribution of human data enhanced the policy’s generalization capability to unseen initial states, enabling successful grasping at different positions or over long distances.
-In the Multi-Object Cluttered Scene, policies trained solely on UMI data demonstrated almost no generalization ability in complex scenarios.
-When factors such as object height or layout changed, policies trained only on UMI data failed, whereas the inclusion of human data gradually enabled success by effectively introducing both the action and observation domains of human data into the robot dataset.
+In the Multi-Object Cluttered Scene, policies trained solely on robot data demonstrated almost no generalization ability in complex scenarios.
+When factors such as object height or layout changed, policies trained only on robot data failed, whereas the inclusion of human data gradually enabled success by effectively introducing both the action and observation domains of human data into the robot dataset.
 Furthermore, the consistently low success rates across all methods in certain tasks highlight important directions for future research: how to further improve the collection precision of human data and how to design efficient algorithms that can better transfer knowledge from human data to dexterous manipulation tasks.
-Appendix C Ethics and Privacy Statement
+Appendix 0.E Ethics and Privacy Statement
 Our work involves large-scale data collection of real-world human manipulation activities. We took a number of measures to ensure responsible data handling, privacy protection, and ethical compliance.